OpenStudy (anonymous):

Find the derivative of f(x) = 8/x at x = -1.

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

do you know the power rule?

OpenStudy (anonymous):

no sorry

OpenStudy (anonymous):

@Peaches15

OpenStudy (anonymous):

here is if this helps f(x) = 8x^(-1) is the same thing as f(x) = 8/x

OpenStudy (anonymous):

oh cool

OpenStudy (anonymous):

now I know the power rule

OpenStudy (anonymous):

so just bring the -1 down and subtract 1 from the exponent of -1

OpenStudy (anonymous):

you make no sense lol

OpenStudy (anonymous):

can you pls show me?

OpenStudy (anonymous):

f(x) = x^6 • 7e^(7x) + e^(7x) • 6x^5 = e^(7x) (7x^6 + 6x^5) hope this help

OpenStudy (anonymous):

yeah afcourse y = 8x^(-1) y'= (-1)8x^(-1-1) y' = -8x^(-2)

OpenStudy (anonymous):

ok what does that do? how does that bring us closer to solving the problem. Pls just list the steps and how to do them for this problem and once I get it once I won't need help again. I will be your 2nd fan and give you medal

OpenStudy (anonymous):

when you have negative exponents you can bring them to the denominator and then they become positive so youll get y' = -8/x^2

OpenStudy (anonymous):

ok but how do i find the gosh darn derivative?

OpenStudy (anonymous):

here plug in your problem into this calculator it shows steps and names the rules that you need to use itll explain it better than i could http://www.derivative-calculator.net/

OpenStudy (anonymous):

y' = -8/x^2 is the derivative

OpenStudy (anonymous):

oho thanks

OpenStudy (anonymous):

see but these are my answers: 4 0 8 -8

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

yeah know you need to plug in the -1 remember from your x = -1 in the question

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

ok thanks got it now